### v3

@@ Appendix D.2  Conditions for Bayes Optimality
$f^{*}_{\pi}({\boldsymbol{x}})({\boldsymbol{\theta}})=\begin{cases}\frac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}&\text{if}\ {\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}}),\\ 0&\text{o.w.}\end{cases}$ (15)
@@
${\mathbb{E}}\big[EU(g,{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(g,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})$ (16)
$=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\Bigg[\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})g({\boldsymbol{x}})({\boldsymbol{\theta}})\Bigg]\pi({\boldsymbol{\theta}})$ (17)
@@
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\Bigg[\max_{{\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}}\Big\{\mathcal{L}({\boldsymbol{\theta}}^{\prime}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}^{\prime})\Big\}\underbrace{\bigg(\sum_{{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})}\frac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}\bigg)}_{=1}\Bigg]$
@@
For alternative utility functions or priors, the optimal rule may not take a convenient form such as the MLE, but can still be computed by solving the linear programming problem of maximizing Bayes expected utility in (17) with respect to the (finitely-many) values of $g({\boldsymbol{x}})({\boldsymbol{\theta}})$, such that $\sum_{\boldsymbol{\theta}}g({\boldsymbol{x}})({\boldsymbol{\theta}})=1$ for each $\boldsymbol{x}$}.
Appendix E  Construction of Credible Sets
We construct credible sets on the joint distribution of potential outcomes in the sample using the posterior belief implied by a uniform prior and the design-based likelihood. A credible set of level $\alpha$ is a set of joint distributions of potential outcomes covering a posterior probability mass of at least $\alpha$. We report “smallest credible sets”, which is the credible set containing the fewest points. We construct these sets with the following algorithm:
@@ Appendix E  Construction of Credible Sets
To report credible sets on the number of always takers, compilers, defiers, or never takers individually, we report the range of values for that type present in the credible set constructed on the joint distribution. Note that the credible sets on individual response types are not independent. For example, the maximal values in the individual credible sets for always takers, compilers, defiers, and never takers may not constitute an element of the credible set on the joint distribution of potential outcomes.
Appendix F  Monty Hall: A Design-Based Model of the Monty Hall Game Can Reveal Evidence to Switch Doors
Suppose you are a contestant on the game show, “Let’s Make a Deal,” with host Monty Hall. He shows you three doors, he tells you that one has a car behind it, and he asks you to choose one. You get the car if you choose the correct door and nothing otherwise. You choose a door. Of the two doors you did not choose, Monty must open one door with no car behind it. Assume that a fair randomizer tells him which remaining door to open if he has two options. Without loss of generality, Monty reveals no car in the left door. Should you keep your door or switch doors?
By well-trodden Bayesian arguments, you should switch doors. Using our proposed visualization of potential outcomes, we illustrate here that the principle of maximum entropy can be used to arrive at the same decision without specifying a Bayesian prior. Wang, Tran, and Abebe (2016) discuss the principle of maximum entropy in the same context. We illustrate it here with our proposed visualization of potential outcomes.
Figure F.2 illustrates the Monty Hall problem using the approach to visualize potential outcomes that we develop to illustrate the design-based likelihood in Figure 1. In Figure 1, the left half of each circle represents the potential outcome in intervention, and the right half of each circle represents the potential outcome in control. The potential outcomes can take on values of “takeup” (orange), “no takeup” (white), or “unobserved” (grey). Here, the two halves of each circle represent the potential outcomes in the two doors that you did not choose: the left half represents the potential outcome in the left remaining door, and the right half represents the potential outcome in the right remaining door. The potential outcomes can take on values of “car” (orange), “no car” (white), or “unobserved” (grey).
Figure F.2: A Design-Based Model of the Monty Hall Game Can Reveal Evidence to Switch Doors
In Figure 1, the left half of the rectangle represents “intervention,” and the right half of the rectangle represents “control.” Here, the design is more complicated because the two possible randomized assignments for Monty are more complicated. The left half of the rectangle represents “reveal left if no car in left, otherwise reveal no car in right” and the right half of the rectangle represents “reveal right if no car in right, otherwise reveal no car in left.”
The data that you observe is “no car in left” and “unobserved in right.” Unlike in Figure 1, you do not observe the randomized assignment (the position of the circle in the rectangle). The two rows depict the two possible joint distributions of potential outcomes that could yield the data. In the first row, in the two doors you did not choose, there is no car in left and no car in right. In the second row, in the two doors you did not choose, there is no car in left, but there is a car in right.
To decide between these two joint distributions of potential outcomes using a design-based maximum likelihood rule and the principle of maximum entropy, you calculate the likelihood in each row, depicted in the last column. In the first row, only one of the two possible randomized assignments for Monty yields the data (no car in left and unobserved in right), so the entropy is 1 and the likelihood is 1/2. In the second row, there are two randomized assignments for Monty that yield the data. (Note that in the right of the rectangle, you can observe the left potential outcome, which is not the case in the model of the experiment in Figure 1.) In the second row, the entropy is 2 and the likelihood is 1. By a design-based maximum likelihood decision rule, the maximum of the likelihood indicates that the car is behind the right remaining door, as depicted in the second row, so you should switch doors. By the principle of maximum entropy, you arrive at the same decision that you would have made under a Bayesian decision rule with a uniform prior.

### v4

@@ Appendix C.2  Conditions for Bayes Optimality
$f^{*}_{\pi}({\boldsymbol{x}})({\boldsymbol{\theta}})=\begin{cases}\dfrac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}&\text{if}\ {\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}}),\\ 0&\text{o.w.}\end{cases}$ (16)
@@
${\mathbb{E}}\big[EU(g,{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(g,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})$
$=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\Bigg[\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})g({\boldsymbol{x}})({\boldsymbol{\theta}})\Bigg]\pi({\boldsymbol{\theta}}).$
@@
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\Bigg[\max_{{\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}}\Big\{\mathcal{L}({\boldsymbol{\theta}}^{\prime}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}^{\prime})\Big\}\underbrace{\bigg(\sum_{{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})}\frac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}\bigg)}_{=1}\Bigg].$
@@
For alternative utility functions or priors, the optimal rule may not take a convenient form such as the MLE, but can still be computed by solving the linear programming problem of maximizing Bayes expected utility in (14) with respect to the (finitely-many) values of $f({\boldsymbol{x}})({\boldsymbol{\theta}})$, such that $\sum_{\boldsymbol{\theta}}f({\boldsymbol{x}})({\boldsymbol{\theta}})=1$ for each $\boldsymbol{x}$}.
Appendix D  Construction of Credible Sets
We construct credible sets on the joint distribution of potential outcomes in the sample using the posterior belief implied by a uniform prior and the design-based likelihood. A credible set of level $\alpha$ is a set of joint distributions of potential outcomes covering a posterior probability mass of at least $\alpha$. We report the smallest credible set, which is the credible set containing the fewest points. We construct these sets with the following algorithm:
@@ Appendix D  Construction of Credible Sets
To report credible sets on the number of always takers, compliers, defiers, or never takers individually, we report the range of values for that type present in the credible set constructed on the joint distribution. Note that the credible sets on individual response types are not independent. For example, the maximal values in the individual credible sets for always takers, compliers, defiers, and never takers may not constitute an element of the credible set on the joint distribution of potential outcomes.
Appendix E  Auxiliary Statistics
Figure E.1: Standard Statistics, Proposed Design-Based Maximum Likelihood Estimates, and Auxiliary Statistics